OpenStudy (anonymous):

can (a^2+a+1) be factored

sam (.sam.):

No

sam (.sam.):

If you try to factor and didn't get any whole number, then the equation cannot be factored even though there is decimal number.

OpenStudy (anonymous):

That's what I thought I just wanted to make sure

OpenStudy (anonymous):

Try the method: \[-b \pm \sqrt{b ^{2}-4ac}/2a=-a \pm \sqrt{a ^{2}-4(a)(1)}/2a=a \pm \sqrt{a ^{2}-4a}/2a\] Don't think it can be factored..

OpenStudy (anonymous):

try quad eq